# Effective Magnetic Switching of THz Signal in Planar Structured Spintronic Emitters

### v2

@@ I Introduction
In the last decade, promising sources of THz radiation based on nanostructures consisting of nano-thick ferromagnetic layers (FM) having boundaries with nano-thick layers of a non-magnetic (usually heavy) metal (NM) have been actively studied and promoted [1, 2, 3, 4, 5]. The minimal effective structure consists of two layers: FM/NM. Numerous studies have shown that such structures can act as broadband THz emitters, comparable in efficiency to semiconductor analogues, ZnTe, semiconductor-based antennas (see for review [6, 1, 4, 7, 3, 8]). Besides they have the sufficient advantage because THz emission can be manipulated by magnetic field in this case [9].
@@ II Methods and Samples
The spintronic terahertz emitters are Co(2)/Pt(2) bilayer films (hereafter, thickness is in nm, order of layers is from substrate) grown on quartz substrates by magnetron sputtering. An external magnetic field of approximately $1kOe$ was applied during growth in order to induce a uniaxial anisotropy in the sample plane. Details of growth technology may be found elsewhere [21]. The $1cm\times 1cm$-size gratings with microns to hundreds of microns period ($4,50,100,300,500,1000\mu m$) and $1/2$ filling factor are then made from these films by optical lithography (Figure 1(a)). We also make a sample with $8\mu m$ period and $3/4$ filling factor ($6\mu m$ stripe width). Samples of periodic spintronic emitters with anisotropy both along the easy axis (E.A.) relative to the stripes and across the stripes were manufactured.
@@ III Experimental results
All samples are verified via the MOKE measurements along and perpendicular to the easy axis. Typical MOKE curves are shown in Figure 2 for the initial film and a grating with $2\mu m$ wire width.
@@
Figure 2: MOKE hysteresis loop along (a) the easy axis (E.A.) and (b) the hard axis (H.A.) for the initial Co(2)/Pt(2) film (solid line) and for the sample with a $4\mu m$ period grating and a $2\mu m$ stripe width (dashed line). The inset shows the orientation of easy axis with respect to the stripe direction.
For micron-sized stripes, the easy axis (E.A.) is along the magnetic field applied during the sample growth regardless of the stripe direction. The magnetization reverses abruptly in $20-40Oe$ if the magnetic field is applied along the easy axis for all the samples. For the magnetic field applied along the hard axis (H.A.) the saturation is reached at approximately $100Oe$.
@@
We perform time dependence of THz field amplitude for the setup when the grating is irradiated by an optical pump from front (sample) and back (substrate) side. The main quantitative difference is the magnitude of signal due to different absorption coefficients for optical and THz signal in a quartz substrate. Note that the results are qualitatively the same both for front and back side. Typical result for irradiation from front side is shown in Figure 1(d) for $1000\mu m$ and $500\mu m$ grating period with $1/2$ filling factor and close geometry.
Figure 3 shows the hysteresis loop for the THz signal when an external magnetic field is applied either along or perpendicular to the easy axis for different grating periods. The hysteresis loop is constructed as follows. The THz signal is measured at a delay position of 0 ps, which corresponds to the peak maximum (as shown in Figure 1(d)). Then, at this position, a magnetic field is applied, and the corresponding THz amplitude is recorded for each value of the magnetic field. Typically, the THz signal polarized in the direction perpendicular to the magnetic field applied to the sample is recorded. We designate this usual signal polarization as X-polarization. In the left column of Figure 3, the hysteresis loops are shown for samples with the E.A. perpendicular to the stripes at grating periods of (a) $1000\mu m$, (b) $50\mu m$, (c) $8\mu m$. In the right column of Figure 3, the hysteresis loops are shown for samples with the E.A. parallel to the stripes at grating periods of (d) $300\mu m$, (e) $100\mu m$, (f) $8\mu m$. The insets at the top of Figure  3 show the orientation of the easy axis (E.A.) and the hard axis (H.A.) relative to the direction of the grating. The black curves correspond to magnetization along the easy axis (E.A.), while the red curves correspond to magnetization along the hard axis (H.A.).
@@
Figure 3: THz hysteresis loop for the magnetic field along (black line) or perpendicular (red line) to easy axis for (a) $1000\mu m$, (b) $50\mu m$, (c) $8\mu m$ grating period and E.A. perpendicular to stripes and (d) $300\mu m$, (e) $100\mu m$, (f) $100\mu m$ grating period and E.A. parallel to stripes. The orientations of easy and hard axis are shown in the insets.
One can straightforwardly see that the signal amplitude in a saturating magnetic field depends on whether the sample is magnetized along or perpendicular to the stripes while the loop shape depends on the direction of easy axis with respect to grating orientation. The decrease in THz signal amplitude after the samples are magnetized along the hard axis and then the magnetic field is turned off is due to the change of the magnetic state: the samples are split into magnetic domains in which the magnetization turns along local easy axis direction. Non-zero remanent magnetization is due to a non-uniform polycrystalline structure of the sample.
@@
To conclude this section, Figure 4 shows the THz signal from the sample with $2\mu m$-wide stripes and a period of $4\mu m$, with the easy axis aligned along the stripes. In Figure 4(a), the time-domain THz signal is shown for different magnetic field orientations: +0.4 kOe and -0.4 kOe along the hard axis (H.A.) and +0.4 kOe and -0.4 kOe along the easy axis (E.A.). For the THz signal measurements along the easy axis (E.A.), the signal was amplified 10 times for better visibility in the plot. The blue line at 0 picoseconds corresponds to the peak maximum, as shown in Figure 4(b). In Figure 4(b), the THz peak amplitude is plotted as a function of the applied magnetic field. It can be seen that there is a drastic 27-fold difference in the signal magnitude upon a 90-degree rotation of the sample.
Figure 4: (a) Time dependence of THz signal in a saturating field of ±$\pm 0.4kOe$ for the sample with $2\mu m$ wide stripes magnetized along hard axis (perpendicular to stripes) and along easy axis (parallel to stripes).(b) THz hysteresis loops and the direction of easy axis and hard axis.
@@ IV Discussion
The results for the THz signal ratio versus the stripe width extracted from the experimental data for different stripe width from $150\mu m$ down to $2\mu$ for both front and back side measurements are shown in Figure 5 (in logarithmic scale).
One can see that the ratio grows as the stripe width becomes smaller, but quite slowly. Similar dependencies are observed for front side and back side measurements.
In order to provide a fit of the experimental data we develop the following theoretical model. For the sample magnetized perpendicularly to the stripes, they are supposed to be infinite. Hence the electric current flows with no charge accumulation. We suppose that the power injected into a unit square of the emitter from magnetic system $p_{i}$ is constant. The electric field emitted per one stripe is then proportional to the electric current through it: $I_{\parallel}\sim\sqrt{\frac{2p_{i}l^{2}h}{\rho}}$, where $\rho$ is resistivity, $l,h$ are length and thickness of the stripe, respectively. For sample magnetization along the stripes a permanent current would split charges with the velocity of light and then stop. Since the current source provided by ISHE is alternating we ought to calculate the high-frequency conductance. It consists of active resistance determined by film conductivity and reactance mainly determined by electrical capacitance of the system. We may roughly estimate it by supposing that the charge is conducted at very thin wires with the diameter equal to the layer thickness $h$ (Figure 6).
@@
Figure 6: A schematic view of stripes cross section. The electric charges accumulate at the boundaries and create an electric field which leads to energy accumulation.
We then calculate the energy power accumulated at such an edge in the electric field of all other edges with alternating charge sign and take into account that the electric field of the charge that lies at the metal side is screened by a factor of $\xi$:
$p_{c}=\left(1-\xi\right)\frac{I_{\perp}^{2}}{2l\omega}Log\left(\frac{L}{w}\right),$ (1)
where $\omega$ is the frequency of electric current $I_{\perp}$, $L$ and $w$ are the characteristic grating width and the stripe width. Taking into account power absorption due to the wire resistance it is then easy to obtain an estimation of the electric current that emits THz signal:
$I_{\perp}\sim\sqrt{\frac{2p_{i}l^{2}h}{\rho+\frac{1-\xi}{\omega}\frac{h}{w}Log\left(\frac{L}{w}\right)}}.$ (2)
It is then straightforward to calculate the THz signal ratio as:
$Ratio=\frac{I_{\parallel}}{I_{\perp}}=\sqrt{1+\frac{h}{L}\frac{1-\xi}{\rho\omega}\frac{L}{w}Log\left(\frac{L}{w}\right)}.$ (3)
@@
The results of experimental data fit with the use of (3) are shown in Figure 5 by solid (dashed) line for front (back) side configuration. Note that we use data for relatively big ($>10\mu m$) stripe width for the fit. The characteristic grid coherence length $L$ is determined from the fit and is equal to $264\mu m$ for front side measurements and $184\mu m$ for back side ones. These values are of the same order; the difference may be attributed to sensitivity of this parameter to small deviations of experimental values caused by noise. One can note that although the real grid size is of the order of $1cm$ we obtain $L$ of the order of THz signal wavelength. This is governed by the fact that equation (3) is obtained in the quasistatic approximation which works well for distances less than the wavelength. The other fitting parameter is the stripe effective resistivity $\rho$. Supposing that the screening parameter $\xi<<1$ we evaluate it from our fit as $\rho=8\cdot 10^{-8}\Omega\cdot m$ for measurements from both sides. Considering the structure as a parallel connection of Pt and Co conductors with the same cross section and resistivity equal to $10.6\cdot 10^{-8}\Omega\cdot m$ and $6.2\cdot 10^{-8}\Omega\cdot m$ respectively [23] we arrive at a very close resistivity equal to $8.4\cdot 10^{-8}\Omega\cdot m$. Thus our rough model fits the experimental data quite well.
@@
It is worth noting that the model (and the experiment) is almost not sensitive to the distance between stripes unless this distance is of the order of value of the stripe width $w$. This is explained by a logarithmic dependence of the energy stored in the accumulated charge on this distance which leads to a small constant correction under the root in (3). Indeed, the point put for $6\mu m$ stripe width in Figure 5 is measured at the $2\mu m$ distance between stripes. Although it is not used for theoretical fit it corresponds to that fit quite well. On the other hand, the ratio value equal to $27$ obtained for the $2\mu m$ stripe width is slightly greater than that predicted by Equation 3 (approximately $15$). This may be attributed to the fact that we do not take into account some other effects such as the circuit inductance in our model. This makes the THz signal ratio even greater.

### v4

@@ I Introduction
In the last decade, promising sources of THz radiation based on nanostructures consisting of nano-thick ferromagnetic layers (FM) having boundaries with nano-thick layers of a non-magnetic (usually heavy) metal (NM) have been actively studied and promoted [1, 2, 3, 4, 5, 6]. The minimal effective structure consists of two layers: FM/NM. Numerous studies have shown that such structures can act as broadband THz emitters, comparable in efficiency to semiconductor analogues, ZnTe, semiconductor-based antennas (see for review [7, 1, 4, 8, 3, 9]). Besides they have the sufficient advantage because THz emission can be manipulated by magnetic field in this case [10].
@@ II Methods and Samples
The spintronic terahertz emitters are Co(2)/Pt(2) bilayer films (hereafter, thickness is in nm, order of layers is from substrate) grown on quartz substrates by magnetron sputtering. An external magnetic field of approximately $1kOe$ was applied during growth in order to induce a uniaxial anisotropy in the sample plane. Details of growth technology may be found elsewhere [22]. The $1cm\times 1cm$-size gratings with microns to hundreds of microns period ($4,50,100,300,500,1000\mu m$) and $1/2$ filling factor are then made from these films by optical lithography (Figure 1(a)). Two series of samples were made. In the first series, the stripes were lithographed along the magnetic anisotropy axis of the original film. In the second series, the stripes were lithographed across the anisotropy axis of the original film. We also make a sample with $8\mu m$ period and $3/4$ filling factor ($6\mu m$ stripe width).
@@ III Experimental results
All samples are verified via the MOKE measurements along and perpendicular to the easy axis. Typical MOKE curves are shown in Figure 2(a,b) for the grating with $8\mu m$ period and $6\mu m$ wire width.
@@
Figure 2: MOKE hysteresis loop (a) along the easy axis (E.A.) and (b) perpendicular to the easy axis (E.A.) for the sample with a $8\mu m$ period grating and a $6\mu m$ stripe width and different orientation of easy axis with respect to the stripe direction, as shown in the inset. (c, d) THz loops for the corresponding geometry of samples and measurement.
Measurements of magnetization loops show that lithography of the lattice does not change the direction of magnetic anisotropy in the sample; it retains the same direction that it had in the original film. Therefore, in the first series of samples (as introduced in Section II), the easy axis of magnetic anisotropy is directed along the lattice stripes, and in the second series of samples, across the lattice. When magnetized along the anisotropy axis, the samples have a rectangular hysteresis loop, the coercivity field is in the range of 20-100 Oe. When magnetized across the easy magnetization axis, the samples demonstrate an inclined hysteresis loop with a residual magnetization in the range of 0.2-0.3 of saturation magnetization and a saturation field in the range of 100-300 Oe. Non-zero remanent magnetization in this case is due to a polycrystalline structure of the sample.
@@
We perform time dependence of THz field amplitude for the setup when the grating is irradiated by an optical pump from front (sample) and back (substrate) side. The main quantitative difference is the magnitude of signal due to different absorption coefficients for optical and THz signal in a quartz substrate. Note that the results are qualitatively the same both for front and back side. Typical result for irradiation from front side is shown in Figure 1(d) for $1000\mu m$ and $500\mu m$ grating period with $1/2$ filling factor and close geometry. The THz hysteresis loops are constructed based on these measurements as follows. The THz signal is measured at a delay position of 0 ps, which corresponds to the peak maximum (as shown in Figure 1(d)). Then, at this position, a magnetic field is applied, and the corresponding THz amplitude is recorded for each value of the magnetic field. Typically, the THz signal polarized in the direction perpendicular to the magnetic field applied to the sample (with the magnetic field along the y-direction) is recorded. We designate this usual signal polarization as X-polarization.
The main experimental results of the article are presented in Figure 2(c,d) and can be briefly formulated as follows. The shape of the dependence of the THz generation amplitude on the applied magnetic field repeats the shape of the magnetization loop of the samples and is determined by the orientation of the field relative to the magnetic axis of anisotropy. Conversely, the maximum generation amplitude observed in saturation depends on the direction of the external magnetic field relative to the orientation of the lattice stripes. An observed discrepancy in hysteresis loop width for MOKE and THz measurements may be attributed to the processes such as sample heating by laser beam [24].
We showed above that the shape of THz loop depends only on the E.A. direction with respect to the applied magnetic field and is almost insensitive to the direction of stripes. Typical THz hysteresis loops for large ($300\mu m$) and small ($4\mu m$) period are shown in Figure 3(a,b) for the case when E.A. is along the stripes. One can note that the THz signal in a saturating magnetic field ($400Oe$) is almost independent of the field direction for grating with large period, whilst demonstrating a strong (approximately 27-fold) decrease upon a 90-degree rotation from the H.A. (perpendicular to E.A.) direction to the E.A. one.
@@
Figure 3: (a,b) THz hysteresis loop for the lattices with 1/2 filling factor and grating period of $300\mu m$ and $4\mu m$, correspondingly. The stripes are along E.A.; the magnetic field is applied along (black line) or perpendicular (red line) to the easy axis of magnetic anisotropy. (c) Dependence of THz amplitude in a saturating magnetic field applied along (black squares) or perpendicular (red circles) to the stripes. The lines are guide for eyes.
The THz peak-to-peak amplitude in a saturating magnetic field applied along the stripes or perpendicular to them for different grating period is shown in Figure 3(c). Note that since the direction of E.A. strongly influences the loop shape leaving the THz signal in a saturating field almost intact Figure 3(c) collects all the data for samples with different directions of magnetization easy axis. This data is now sorted by the direction of the magnetic field with respect to the stripes (a $2/3$ scaling factor is used for the sample with $6\mu m$ stripe width and $8\mu m$ period). It is evident that if the magnetic field is oriented across the gratings, the magnitude of the THz signal in saturation weakly depends on the grating period $T$. A minor decrease for small periods may be attributed to the fact that sufficient power is emitted into side diffraction maxima which are not collected by the parabolic mirror in our setup. Conversely, if the magnetizing field is oriented along the gratings, then a decrease in the grating period leads to a significant decrease in the THz signal. In other words, for gratings with large periods ($1000,500,300\,nm$), turning the saturating field relative to the gratings does not lead to a change in the amplitude of THz generation. Meanwhile, for gratings with a small period ($4nm$), a 90-degree turn of the saturating field from perpendicular to parallel direction relative to the grating stripes leads to a drastic 27-fold decrease in the generation amplitude.
Changes in the amplitude of THz generation during rotation of the magnetic field relative to the grating depending on the grating period are summarized in Figure 4 as the ratio of corresponding THz signals (in logarithmic scale).
One can see that the ratio decreases as the stripe width increases, but quite slowly. Similar dependencies are observed for front side and back side measurements. We include the results obtained from the structure with a $8~{}\mu m$ period (Figure 2(c,d)) and 3/4 filling factor into Figure 4. Contrary to our expectations, the absolute THz signal from these structures is not greater than from the structures with 1/2 filling factor (Figure 3(c)). However the ratio corresponds to theoretical fit (described below) based on their stripe width quite well.
Spectral properties of the THz signal obtained from the stripes with different representative periods in a saturating magnetic field applied perpendicular to the stripes are shown in Figure 5. A weak shift of the radiation bandwidth to the high frequency range is observed for small grating period. This shift is absent for $100\mu m$ period and is roughly $10\%$ (or $0.07THz$) for $4\mu m$ period. This may be explained by the fact that very small size of stripes as emitters compared to wavelength becomes important. This slightly suppresses emission for long wavelength while for small wavelength (high frequency) the grating still works as a sum of rather big rectangles.
Figure 5: Normalized spectra of THz signal for different stripe periods. A saturating magnetic field of $400Oe$ is applied perpendicular to the stripes.
@@ IV Discussion
We develop a simple theoretical model in order to explain the experimental data. The model is described in detail in Supplementary Materials; here we briefly point out its main features and pass to the final result for the THz signal ratio which is then used to fit the experimental data.
For the sample magnetized perpendicularly to the stripes, the electric current provided by ISHE flows along the stripes which are supposed to be very long (we neglect charge accumulation at the stripes’ ends). We suppose that the power injected into a unit square of the emitter from magnetic system $p_{i}$ is constant. The electric field emitted by such system is then proportional to the electric current density averaged over the square:
$j_{\parallel}\right>\sim\sqrt{\frac{p_{i}}{2\rho h}}.$ (1)
where $\rho$ is resistivity, $h$ is the thickness of the stripe.
A different case occurs if the sample is magnetized along the stripes: the current caused by ISHE flows perpendicular to the stripes which leads to charge accumulation (Figure 6).
@@
Figure 6: A schematic view of stripes cross section. The power $p_{i}$ is injected from a Co ferromagnet (sand color) to a Pt heavy metal (purple color) of a bilayer structure lying on an optically transparent quartz substrate (blue). The electric charges accumulate at the Pt boundaries and create an electric field which leads to energy accumulation.
In order to determine the current in this case we ought to calculate the high-frequency conductance taking into account active resistance determined by film conductivity and reactance mainly determined by electrical capacitance of the system. We may roughly estimate it by supposing that the charge is conducted at very thin wires with the diameter equal to the layer thickness $h$. We then calculate the energy power accumulated at such a charged edge in the system of stripes and taking into account power absorption due to the wire resistance obtain an estimation of the electric current electric current density averaged over the square in this case as:
$j_{\perp}\right>\sim\sqrt{\frac{p_{i}}{2h\left(\rho+\frac{2}{\omega}\frac{h}{w}Log\left(\frac{L}{w}\right)\right)}},$ (2)
where $L$ and $w$ are the characteristic grating width and the stripe width, $\omega$ is the frequency of electric current. The THz signal ratio follows from (1) and (2):
[…]
@@
The results of experimental data fit with the use of (3) are shown in Figure 4 by solid (dashed) line for front (back) side configuration. Note that we use data for relatively big ($>10\mu m$) stripe width for the fit. The characteristic grid coherence length $L$ is determined from the fit and is equal to $264\mu m$ for front side measurements and $184\mu m$ for back side ones. These values are of the same order; the difference may be attributed to sensitivity of this parameter to small deviations of experimental values caused by noise. One can note that although the real grid size is of the order of $1cm$ we obtain $L$ of the order of THz signal wavelength. This is governed by the fact that equation (3) is obtained in the quasistatic approximation which works well for distances less than the wavelength. The other fitting parameter is the stripe effective resistivity $\rho$. We evaluate it from our fit as $\rho=1.6\cdot 10^{-7}\Omega\cdot m$ which is the same for measurements from both sides. Considering the structure as a parallel connection of Pt and Co conductors with the same cross section and resistivity equal to $10.6\cdot 10^{-8}\Omega\cdot m$ and $6.2\cdot 10^{-8}\Omega\cdot m$ respectively [25] we arrive at the structure resistivity equal to $8.4\cdot 10^{-8}\Omega\cdot m$ which is approximately two times smaller than obtained from our fit. This is of the same order of value; the difference is explained by the fact that thin films usually have greater resistivity than bulk material.
@@
It is worth noting that the model (and the experiment) is almost not sensitive to the distance between stripes unless this distance is of the order of value of the stripe width $w$. This is explained by a logarithmic dependence of the energy stored in the accumulated charge on this distance which leads to a small constant correction under the root in (3). Indeed, the point put for $6\mu m$ stripe width in Figure 4 is measured for the structure with $8\mu m$ period and 3/4 filling factor. It corresponds to the fit quite well. On the other hand, the ratio value equal to $27$ obtained for the $2\mu m$ stripe width is slightly greater than that predicted by Equation 3 (approximately $15$) which may be attributed to the fact that we do not take into account some effects such as the circuit inductance in our rough model. This makes the THz signal ratio even greater.